Vacuum work is completed. The TP2 and TP3 interlocks have been overhauled as proposed in ELOG 15499 and seem to be performing reliably. We're now back in the nominal system state, with TP2 again backing for TP1 and TP3 pumping the annuli. I'll post the full implementation details in the morning.
I did not get to setting up the new UPS units. That will have to be scheduled for another day.
The vac system is going down now for planned repairs [ELOG 15499]. It will likely take most of the day. Will advise when it's back up.